OpenStudy (anonymous):

Find an equation of the tangent line to the graph at the given point. (x + 2)^2 + (y − 3)^2 = 37, (−1, −3) i got -(x+2)/(y-3) and when i plug in for x and y i got (y+3)=1/6(x+1) but it keeps saying my answer is wrong what am i doing wrong

OpenStudy (anonymous):

what is dy/dx?

OpenStudy (anonymous):

-(x+2)/(y-3)

OpenStudy (anonymous):

thats what i got for the dy/dx and then when i plug in points i got (y+3)=1/6(x+1) but it said thats wrong

OpenStudy (anonymous):

so what is dy/dx, meaning what does it represent?

OpenStudy (anonymous):

you use dy/dx when you take the derivative of y

OpenStudy (anonymous):

i know how to take derivatives and all and i know the answer i got was correct for the dy/dx because i check on multiple sights like alpha and all and came out with same answer i got so i have to somehow made a mistake plugging in

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

it's the slope... you have to use the point to get the slope at that point. then use the slope and the point in the point-slope form.

OpenStudy (anonymous):

you may have taken the derivative correctly but your interpretation and how you are using it is not correct.

OpenStudy (anonymous):

so how is (y+3)=1/6(x+1) not right because i been using the same form on every other problem i been doing this way and that form was right

OpenStudy (anonymous):

im confused lol

OpenStudy (anonymous):

slope at (-1, -3)

OpenStudy (anonymous):

\[\frac{ dy }{ dy } = -\frac{ x+2 }{ y-3 } \Rightarrow -\frac{ \left( -1 \right)+2 }{ \left( -3 \right)-3 } =-\frac{ 1 }{ -6 } =\frac{ 1 }{ 6 }\] slope intercept form: y-(-1)=1/6(x-(-3)) => y+1 = 1/6 (x+3)

OpenStudy (anonymous):

ooops y+3 = 1/6(x+1)

OpenStudy (anonymous):

yeah, i see your point... let me check dy/dx

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

2(x+2)+2(y-3)y'=0 y' = -(x+2)/(y-3)

OpenStudy (anonymous):

yep not sure of the issue. sorry i didn't pay closer attention at the start.

OpenStudy (anonymous):

wait i forgot it does say circle do i need to write it in some form to match that

OpenStudy (anonymous):

could it be it thinks the (x+1) is on the bottom? how you are typing it in? how do you do fractions in your program?

OpenStudy (anonymous):

do they want slope-intercept form? y = mx+b

OpenStudy (anonymous):

im not sure because the fact its suppose to take both forms

OpenStudy (anonymous):

try y = x/6 - 17/6 or y = (x-17)/6

OpenStudy (anonymous):

it took y=(x-17)/6 how did you get that

OpenStudy (anonymous):

never mind i saw how you got it thank you